OpenStudy (anonymous):

Write as a sum or difference of logarithms with no unnecesssry exponents left. ln ((5x^4)(2x-1)^7) / (([square root] x^2 +1)(3x+2)^19)

OpenStudy (anonymous):

\[\ln 5x^4 (2x-1)^7 \div \sqrt{x^2 +1}(3x+2)^19\]

OpenStudy (anonymous):

except the last exponent is 19

OpenStudy (campbell_st):

for the numerator you will need to apply these rules. \[\log(ab) = \log(a) + \log(b)\] and the index rule is \[\log(x^a) = a \times \log(x)..... or....alog(x)\] for the denominator \[\sqrt{(x^2 +1)(3x + 2)^19} = \sqrt{(x^2 +1) \times (3x + 2)^{18} \times (3x + 2)} = \sqrt{(3x +2)^{18}} \times \sqrt{(x^2 +1)(3x + 2)}\] you need to finish simplifying the denominator